OpenStudy (anonymous):

find the minimum y-value on the graph of y=f(x) for f(x)=7x^2+7x-4

OpenStudy (anonymous):

I have gotten as far as \[7(-\frac{ 1 }{ 2 })^2+7(-\frac{ 1 }{ 2 })-4\] then I get stuck

OpenStudy (anonymous):

I know that I have to evaluate \[f(-\frac{ 1 }{ 2 }) \] but I am not sure how to do so

OpenStudy (anonymous):

I have gotten the x-value now I just need the y-value

OpenStudy (anonymous):

I know the steps on how to do this but for some reason this one has me stuck

OpenStudy (psymon):

Well, what is (-1/2)^2?

OpenStudy (anonymous):

wouldnt it be -1

OpenStudy (psymon):

Thats adding them, we want to square it. So (-1/2)*(-1/2)

OpenStudy (anonymous):

I dont know then

OpenStudy (anonymous):

-1/4

OpenStudy (psymon):

Close. you have to multiply both negatives, too. 2 negatives make a positive.

OpenStudy (anonymous):

oh yeah ok so 1/4

OpenStudy (psymon):

Yep. So now 7(-1/2)?

OpenStudy (anonymous):

3 1/2?

OpenStudy (anonymous):

-7/2

OpenStudy (psymon):

Yep, exactly. So you need to be able to do 7/4 - 7/2 - 4.

OpenStudy (anonymous):

and how do I do that

OpenStudy (anonymous):

would I make the 4 a 4/1

OpenStudy (psymon):

Well, you would need a common denominator for all of them.

OpenStudy (anonymous):

\[\frac{ 7 }{ 4 }-\frac{ 7 }{ 2 }-\frac{ 4 }{ 1 }\]

OpenStudy (psymon):

Thats a start. So what would your common denominator be?

OpenStudy (anonymous):

common denominator would be 4

OpenStudy (psymon):

Right. So now do what you need to do to make all of those fractions have a denominator of 4 and then subtract : )

OpenStudy (anonymous):

\[\frac{ 7 }{4? }-\frac{ 14 }{ 4 }-\frac{ 16 }{ 4 }\]

OpenStudy (anonymous):

sorry didnt mean to leave the ?

OpenStudy (psymon):

Looks fine to me. Now just subtract.

OpenStudy (anonymous):

so -23/4

OpenStudy (psymon):

Yep, that would be your y-value :3

OpenStudy (anonymous):

ok thanks, I know this stuff but for some reason that one puzzled me

OpenStudy (psymon):

No worries. You get better and better at them :3 Good luck.